Determine the x-intercepts of a cubic function y= x2 - 2x2- 5x + 6


1
Expert's answer
2021-03-16T08:40:58-0400

y = x³ - 2x² - 5x + 6


"y = 1² - 2(1)² - 5(1) + 6\\

y = 1 -2 -5 +6\\

y = 0\\"


"\therefore" x= 1 is an intercept on the x axis


"\dfrac{x³ - 2x² - 5x + 6}{x-1 } = x² - x - 6"


"x\u00b2 - x - 6 = (x -3)(x+2)"


"\therefore" the x - intercepts of the cubic function are -2, 1 and 3